# HEST-1k: A Dataset for Spatial Transcriptomics and Histology Image Analysis

### v1

@@ F.2 Dataset composition
  • What are the instances? The modalities used in this study are histopathology whole-slide images, gene expression data, and derivatives of these two modalities, such as nuclear segmentation maps.
@@
  • What data does each instance consist of? Imaging data consists of Generic TIFF objects stored in pyramidal format, and gene expression data consists of scanpy objects. Derivatives are stored in JSON and Hierarchical Data Format (HDF) files.
@@
  • Is everything included or does the data rely on external resources? (e.g., websites, tweets, datasets) If external resources, a) are there guarantees that they will exist, and remain constant, over time; b) is there an official archival version. Are there licenses, fees or rights associated with any of the data? We provide all data as part of HEST-1k release. In addition, a link to the original data is provided in the metadata. Each sample is associated with a license as provided by the original publication, where we ensured that the reported license allowed for distributing and creating derivatives of the data. Should the author of the original publication would like to remove their samples from HEST-1k, we will update the data promptly.
@@
  • What experiments were initially run on this dataset? Have a summary of those results and, if available, provide the link to a paper with more information here. At the time of submission, all experiments run with HEST-1k are part of this study. The reader can refer to the main text for a thorough description of all experiments (see HEST-Benchmark, HEST for discovery, HEST for multimodal representation learning).
@@ F.3 Data collection process
  • Who was involved in the data collection process? All authors of the present study were involved in the data collection, inspection, and curation. The reader can refer to the original publication to understand how the data were orginally acquired.
  • Over what time-frame was the data collected? Does the collection time-frame match the creation time-frame? The original data comprise publications from 2016 to 2024. As the dataset grows, more recent publications might be included in HEST-1k.
@@
  • Is there information missing from the dataset and why? (this does not include intentionally dropped instances; it might include, e.g., redacted text, and withheld documents) Is this data missing because it was unavailable? Original publications include some missing information, such as the alignment file between the slide and the expression profile. We developed computational tools to minimize missing information to reach near-complete metadata.
  • Are there any known errors, sources of noise, or redundancies in the data? All whole-slide images have been manually inspected. The quality from one sample to another varies significantly, for instance, due to poor staining, compression artifact, lower resolution, etc. Gene expression data are inherently noisy. Users can decide to apply post-hoc normalization methods to reduce noise.
@@ F.4 Data preprocessing
  • What preprocessing/cleaning was done? All whole-slide images were converted into a pyramidal TIFF objects with re-estimated pixel resolution. All alignment files have been manually inspected and included if missing. All gene expression data have been transformed into scanpy objects following the same process.
  • Was the “raw” data saved in addition to the preprocessed/cleaned data? (e.g., to support unanticipated future uses) Raw data are downloaded, but not publicly shared. In the case of public samples, users can re-download them using the metadata provided as part of the dataset release.
@@ F.5 Dataset distribution
  • How is the dataset distributed? HEST-1k is distributed using internal storage. We have several internal copies to ensure data preservation over the years.
  • When will the dataset be released/first distributed? The dataset will be made public with the camera-ready version of this publication. This study serves as the canonical reference.
@@
  • Are there any fees or access/export restrictions? No access/export restrictions unless they violates the terms of the above-mentioned license (CC BY-NC-SA 4.0 Deed).
@@
  • Will the dataset be updated? How often and by whom? How will updates/revisions be documented and communicated (e.g., mailing list, GitHub)? Is there an erratum? The dataset might evolve as additional samples become publicly available. Dataset versioning will be put in place. A website will clearly state what the latest version of the dataset contains and keep track of the changes in each version.
  • If the dataset becomes obsolete how will this be communicated? The corresponding website will state if the dataset becomes obsolete.
@@
  • If others want to extend/augment/build on this dataset, is there a mechanism for them to do so? If so, is there a process for tracking/assessing the quality of those contributions. What is the process for communicating/distributing these contributions to users? Users are welcome to contact us if they would like to provide additional data that meet the standards of the data contained in HEST-1k. We do not have a dedicated system to communicate these contributions. Newly added data will be tracked in the versioning.
@@ F.6 Legal and ethical considerations
  • If it relates to people, could this dataset expose people to harm or legal action? (e.g., financial social or otherwise) What was done to mitigate or reduce the potential for harm? No, patients cannot be linked to the corresponding histology and gene expression profile.
@@ Appendix G Author statement
Table A13: HEST-Benchmark results. Ridge regression evaluation on public state-of-the-art patch encoders. Model performance measured with Pearson correlation. We report the mean $\pm$ standard deviation computed over all folds (or patients). Best is bold, second best is underlined.
ResNet50 KimiaNet Ciga CTransPath Remedis Phikon PLIP UNI CONCH Gigapath
IDC 0.446 0.455 0.425 0.499 0.496 0.526 0.450 $\bm{0.544}$ 0.541 $\underline{0.542}$
±$\pm 0.04$$\pm 0.03$$\pm 0.04$$\pm 0.06$$\pm 0.07$$\pm 0.08$$\pm 0.04$$\pm 0.07$$\pm 0.07$$\pm 0.09$
PRAD 0.276 0.323 0.258 $0.290$ 0.264 0.249 0.262 0.257 $\bm{0.353}$ 0.301
±$\pm 0.06$$\pm 0.04$$\pm 0.06$$\pm 0.07$$\pm 0.08$$\pm 0.13$$\pm 0.02$$\pm 0.08$$\pm 0.01$$\pm 0.03$
PAAD 0.343 0.374 0.364 $\underline{0.383}$ 0.288 0.360 0.358 0.377 $\bm{0.424}$ 0.323
±$\pm 0.07$$\pm 0.08$$\pm 0.07$$\pm 0.07$$\pm 0.05$$\pm 0.07$$\pm 0.06$$\pm 0.06$$\pm 0.04$$\pm 0.00$
SKCM $\underline{0.413}$ 0.357 0.381 0.404 0.412 0.369 0.337 0.343 $\bm{0.546}$ 0.190
±$\pm 0.08$$\pm 0.05$$\pm 0.06$$\pm 0.07$$\pm 0.04$$\pm 0.11$$\pm 0.08$$\pm 0.06$$\pm 0.05$$\pm 0.05$
COAD $\underline{0.116}$ 0.103 0.093 0.113 0.075 0.114 0.108 0.110 $\bm{0.135}$ 0.085
±$\pm 0.05$$\pm 0.04$$\pm 0.02$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.05$$\pm 0.06$$\pm 0.06$$\pm 0.03$
READ 0.067 0.058 0.081 0.090 0.077 0.114 0.084 0.124 $\bm{0.154}$ $\underline{0.125}$
±$\pm 0.06$ 0.05 ±$\pm 0.05$$\pm 0.08$$\pm 0.07$$\pm 0.07$$\pm 0.06$$\pm 0.04$$\pm 0.04$$\pm 0.05$
ccRCC 0.210 0.190 0.180 0.214 0.219 $\bm{0.253}$ 0.174 $\underline{0.240}$ 0.221 0.226
±$\pm 0.05$$\pm 0.06$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.05$
HCC $\underline{0.039}$ 0.031 0.027 0.024 0.016 0.038 0.032 0.017 $\bm{0.046}$ 0.008
±$\pm 0.02$$\pm 0.01$$\pm 0.02$$\pm 0.02$$\pm 8.71$$\pm 0.01$$\pm 0.03$$\pm 0.00$$\pm 0.01$$\pm 0.00$
LUAD 0.400 0.419 0.417 0.403 0.311 $\underline{0.422}$ 0.406 0.371 $\bm{0.493}$ 0.297
±$\pm 0.06$$\pm 0.07$$\pm 0.04$$\pm 0.07$$\pm 0.04$$\pm 0.06$$\pm 0.01$$\pm 0.11$$\pm 0.03$$\pm 0.08$
LYM-IDC 0.203 0.227 0.218 0.209 0.167 0.215 0.215 0.224 $\bm{0.247}$ 0.191
±$\pm 0.05$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.04$$\pm 0.02$
Average 0.251 0.253 0.244 0.263 0.232 0.266 0.243 0.260 $\bm{0.316}$ 0.229

### v2

@@ F.2 Dataset composition
  • What are the instances? The modalities used in this study are histopathology whole-slide images, gene expression data, and derivatives of these two modalities, such as nuclear segmentation and classification maps.
@@
  • What data does each instance consist of? Imaging data consists of Generic TIFF objects stored in a pyramidal format, and gene expression data consists of scanpy objects. Derivatives are stored in JSON files, parquet files, and Hierarchical Data Format (HDF) files.
@@
  • Is everything included or does the data rely on external resources? (e.g., websites, tweets, datasets) If external resources, a) are there guarantees that they will exist, and remain constant, over time; b) is there an official archival version. Are there licenses, fees or rights associated with any of the data? We provide all data as part of the HEST-1k release. In addition, a link to the original data is provided in the metadata. Each sample is associated with a license as provided by the original publication, where we ensured that the reported license allowed for distributing and creating derivatives of the data.
@@
  • What experiments were initially run on this dataset? Have a summary of those results and, if available, provide the link to a paper with more information here. All experiments run with HEST-1k are described in this study. The reader can refer to the main text for a thorough description of all experiments (see HEST-Benchmark, HEST for biomarker exploration, HEST for multimodal representation learning).
@@ F.3 Data collection process
  • Who was involved in the data collection process? All authors of the present study were involved in the data collection, inspection, and curation. The reader can refer to the original publication to understand how the data were originally acquired.
  • Over what time frame was the data collected? Does the collection time frame match the creation time frame? The original data comprise publications from 2016 to 2024. As the dataset grows, more recent publications might be included in HEST-1k.
@@
  • Is there information missing from the dataset and why? (this does not include intentionally dropped instances; it might include, e.g., redacted text, and withheld documents) Is this data missing because it was unavailable? Original publications may include some missing information, such as the alignment file between the slide and the expression profile. We developed computational tools to minimize missing information and reach near-complete metadata.
  • Are there any known errors, sources of noise, or redundancies in the data? All whole-slide images have been manually inspected. The quality from one sample to another varies significantly, for instance, due to poor staining, compression artifact, lower resolution, etc. Gene expression data are inherently noisy. Users can decide to apply post-hoc normalization methods to reduce noise, e.g., stain normalization on the imaging side or batch effect mitigation on the transcriptomics side.
@@ F.4 Data preprocessing
  • What preprocessing/cleaning was done? All whole-slide images were converted into pyramidal TIFF objects with re-estimated pixel resolution. All alignment files have been manually inspected and included if missing. All gene expression data have been transformed into scanpy objects following the same process.
  • Was the “raw” data saved in addition to the preprocessed/cleaned data? (e.g., to support unanticipated future uses) Raw data are downloaded but not publicly shared. In the case of public samples, users can re-download them using the metadata provided as part of the dataset release.
@@ F.5 Dataset distribution
  • How is the dataset distributed? HEST-1k is distributed using HuggingFace Datasets.
  • When will the dataset be released/first distributed? The dataset is public and can be accessed through the HuggingFace Datasets interface.
@@
  • Are there any fees or access/export restrictions? No access/export restrictions unless they violate the terms of the above-mentioned license (CC BY-NC-SA 4.0 Deed).
@@
  • Will the dataset be updated? How often and by whom? How will updates/revisions be documented and communicated (e.g., mailing list, GitHub)? Is there an erratum? The dataset might evolve as additional samples become publicly available. Dataset versioning will be put in place.
  • If the dataset becomes obsolete how will this be communicated? The GitHub README will be updated.
@@
  • If others want to extend/augment/build on this dataset, is there a mechanism for them to do so? If so, is there a process for tracking/assessing the quality of those contributions. What is the process for communicating/distributing these contributions to users? Users are welcome to contact us if they would like to provide additional data that meets our standards. We do not have a dedicated system to communicate these contributions. Newly added data will be tracked in the versioning.
@@ F.6 Legal and ethical considerations
  • If it relates to people, could this dataset expose people to harm or legal action? (e.g., financial, social or otherwise) What was done to mitigate or reduce the potential for harm? No, patients cannot be linked to the corresponding histology and gene expression profile.
@@ Appendix G Author statement
Table A13: HEST-Benchmark evaluated using Ridge regression. Model performance measured with Pearson correlation. Best is bold, second best is underlined.
IDC PRAD PAAD SKCM COAD READ ccRCC LUAD LYMPH IDC Average
REMEDIS 0.4936 0.2632 0.2881 0.4117 0.151 0.0776 0.2201 0.3114 0.1694 0.2651
$\pm$ 0.0725 $\pm$ 0.0821 $\pm$ 0.0544 $\pm$ 0.0384 $\pm$ 0.0147 $\pm$ 0.0684 $\pm$ 0.0418 $\pm$ 0.0432 $\pm$ 0.0365
GigaPath 0.532 0.3035 0.3172 0.2231 0.163 0.1236 0.2172 0.3144 0.1925 0.2652
$\pm$ 0.0812 $\pm$ 0.0279 $\pm$ 0.0165 $\pm$ 0.0071 $\pm$ 0.041 $\pm$ 0.0379 $\pm$ 0.0479 $\pm$ 0.0871 $\pm$ 0.0304
UNIv1.5 0.5657 0.3065 0.3004 0.258 0.1982 0.1077 0.2023 0.3084 0.1998 0.2719
$\pm$ 0.0866 $\pm$ 0.02 $\pm$ 0.0199 $\pm$ 0.0514 $\pm$ 0.0262 $\pm$ 0.0222 $\pm$ 0.0595 $\pm$ 0.0747 $\pm$ 0.0129
H-Optimus-0 0.5789 0.2561 0.3367 0.2778 0.1605 0.1228 0.2342 0.3143 0.1976 0.2754
$\pm$ 0.0899 $\pm$ 0.0003 $\pm$ 0.0428 $\pm$ 0.0048 $\pm$ 0.0522 $\pm$ 0.0309 $\pm$ 0.0373 $\pm$ 0.083 $\pm$ 0.0253
Virchow2 0.5666 0.2972 0.2718 0.303 0.1814 0.1208 0.2257 0.3017 0.2172 0.2762
$\pm$ 0.0848 $\pm$ 0.037 $\pm$ 0.0387 $\pm$ 0.0184 $\pm$ 0.0326 $\pm$ 0.0526 $\pm$ 0.0433 $\pm$ 0.1199 $\pm$ 0.019
Virchow 0.5583 0.2744 0.3361 0.3389 0.1825 0.0955 0.2375 0.2897 0.2081 0.2801
$\pm$ 0.0876 $\pm$ 0.0019 $\pm$ 0.037 $\pm$ 0.0063 $\pm$ 0.0369 $\pm$ 0.0527 $\pm$ 0.0371 $\pm$ 0.0785 $\pm$ 0.0234
ResNet50 0.4453 0.2753 0.3432 0.413 0.2009 0.0669 0.2103 0.4001 0.203 0.2842
$\pm$ 0.0377 $\pm$ 0.0622 $\pm$ 0.0654 $\pm$ 0.0814 $\pm$ 0.061 $\pm$ 0.0646 $\pm$ 0.0548 $\pm$ 0.0637 $\pm$ 0.0536
CTransPath 0.4996 0.2895 0.3826 0.4038 0.1751 0.0909 0.2139 0.4026 0.2089 0.2963
$\pm$ 0.0594 $\pm$ 0.0724 $\pm$ 0.066 $\pm$ 0.065 $\pm$ 0.0423 $\pm$ 0.0808 $\pm$ 0.0438 $\pm$ 0.07 $\pm$ 0.0367
Phikon 0.5259 0.2493 0.3594 0.3684 0.1697 0.1136 0.253 0.4224 0.2151 0.2974
$\pm$ 0.0791 $\pm$ 0.1264 $\pm$ 0.0707 $\pm$ 0.1061 $\pm$ 0.0562 $\pm$ 0.0749 $\pm$ 0.0483 $\pm$ 0.0579 $\pm$ 0.0416
UNI 0.563 0.257 0.3768 0.3433 0.1839 0.1239 0.2395 0.3714 0.2236 0.2981
$\pm$ 0.0771 $\pm$ 0.0819 $\pm$ 0.0555 $\pm$ 0.0556 $\pm$ 0.0509 $\pm$ 0.0434 $\pm$ 0.0557 $\pm$ 0.1098 $\pm$ 0.0289
CONCH 0.528 0.3604 0.4224 0.5079 0.2467 0.1443 0.2356 0.4957 0.2462 0.3541
$\pm$ 0.0794 $\pm$ 0.0135 $\pm$ 0.0773 $\pm$ 0.0281 $\pm$ 0.0045 $\pm$ 0.0455 $\pm$ 0.0387 $\pm$ 0.0203 $\pm$ 0.0349
Table A14: HEST-Benchmark evaluated using XGBoost regression. Model performance measured with Pearson correlation. Best is bold, second best is underlined.
IDC PRAD PAAD SKCM COAD READ ccRCC LUAD LYMPH IDC Average
ResNet50 (IN) 0.4646 0.3433 0.4017 0.4707 0.2892 0.0586 0.181 0.4967 0.2284 0.326
$\pm$ 0.0353 $\pm$ 0.0168 $\pm$ 0.0648 $\pm$ 0.0834 $\pm$ 0.0115 $\pm$ 0.069 $\pm$ 0.0502 $\pm$ 0.01 $\pm$ 0.0511
CTransPath 0.4738 0.3514 0.4257 0.5304 0.2921 0.0996 0.2026 0.5225 0.234 0.348
$\pm$ 0.0394 $\pm$ 0.0032 $\pm$ 0.0701 $\pm$ 0.073 $\pm$ 0.0018 $\pm$ 0.0766 $\pm$ 0.0387 $\pm$ 0.0063 $\pm$ 0.0613
Phikon 0.4704 0.3943 0.3988 0.5323 0.277 0.1451 0.213 0.542 0.2443 0.3575
$\pm$ 0.0672 $\pm$ 0.0123 $\pm$ 0.0598 $\pm$ 0.0607 $\pm$ 0.0098 $\pm$ 0.0851 $\pm$ 0.0362 $\pm$ 0.017 $\pm$ 0.0632
GigaPath 0.5222 0.3749 0.4415 0.5297 0.2876 0.1609 0.2207 0.5506 0.2464 0.3705
$\pm$ 0.0641 $\pm$ 0.0103 $\pm$ 0.058 $\pm$ 0.0376 $\pm$ 0.0039 $\pm$ 0.0777 $\pm$ 0.0402 $\pm$ 0.0108 $\pm$ 0.0526
CONCH 0.5175 0.3784 0.4428 0.5766 0.3215 0.1431 0.1738 0.5581 0.2554 0.3742
$\pm$ 0.0602 $\pm$ 0.0124 $\pm$ 0.0657 $\pm$ 0.0519 $\pm$ 0.0062 $\pm$ 0.0665 $\pm$ 0.0544 $\pm$ 0.0081 $\pm$ 0.0605
REMEDIS 0.5116 0.3526 0.4621 0.5885 0.319 0.1129 0.2303 0.562 0.2521 0.3768
$\pm$ 0.0594 $\pm$ 0.0073 $\pm$ 0.0555 $\pm$ 0.0253 $\pm$ 0.0101 $\pm$ 0.0846 $\pm$ 0.0393 $\pm$ 0.0057 $\pm$ 0.0601
Virchow2 0.5378 0.3772 0.4237 0.5565 0.281 0.1779 0.2428 0.5641 0.2582 0.3799
$\pm$ 0.0685 $\pm$ 0.007 $\pm$ 0.0525 $\pm$ 0.0152 $\pm$ 0.0162 $\pm$ 0.077 $\pm$ 0.0361 $\pm$ 0.0069 $\pm$ 0.0504
UNI 0.538 0.3513 0.451 0.6089 0.2921 0.1679 0.235 0.5357 0.2456 0.3806
$\pm$ 0.0603 $\pm$ 0.0162 $\pm$ 0.0587 $\pm$ 0.0294 $\pm$ 0.0191 $\pm$ 0.0641 $\pm$ 0.0381 $\pm$ 0.0057 $\pm$ 0.05
Virchow 0.5309 0.3447 0.4448 0.6089 0.3275 0.1419 0.2307 0.5643 0.2617 0.3839
$\pm$ 0.0764 $\pm$ 0.0117 $\pm$ 0.0501 $\pm$ 0.0165 $\pm$ 0.0254 $\pm$ 0.0669 $\pm$ 0.0336 $\pm$ 0.0091 $\pm$ 0.0537
UNIv1.5 0.555 0.3654 0.434 0.6025 0.336 0.1742 0.2166 0.5634 0.2515 0.3887
$\pm$ 0.0763 $\pm$ 0.0098 $\pm$ 0.0568 $\pm$ 0.0385 $\pm$ 0.0179 $\pm$ 0.0568 $\pm$ 0.0337 $\pm$ 0.0054 $\pm$ 0.0434
H-Optimus-0 0.5564 0.3829 0.4445 0.6502 0.2922 0.1731 0.2402 0.5654 0.2555 0.3956
$\pm$ 0.0777 $\pm$ 0.0049 $\pm$ 0.0563 $\pm$ 0.0326 $\pm$ 0.0063 $\pm$ 0.0777 $\pm$ 0.0348 $\pm$ 0.0084 $\pm$ 0.0522
[…]